OpenStudy (anonymous):

find the inverse of h(x) = 3x+6/5

OpenStudy (anonymous):

@luisrock2008 now i had told you how to get it solve it on your own

OpenStudy (anonymous):

yeah but this is a fraction

OpenStudy (anonymous):

y=h(x)=3x+6/5 x=3y+6/5 <-- it doesn't matter if its a fraction..

OpenStudy (anonymous):

take the fraction to the other side x-6/5=3y (now divide 3 on both sides) x-(6/5)/3=y

OpenStudy (anonymous):

oh, thx

OpenStudy (anonymous):

yw..